[ ESO ] La Silla

EUROPEAN SOUTHERN OBSERVATORY
SciOp

SofI Status News


News Items

  • An intervention was carried out at the beginning of July 2006 and the grism wheel problem was solved. Now SofI is in normal operation! Thanks to all involved in the repairs (July 2006)!
  • SofI underwent an intervention during the first half of Nov 2005 to fix the Slit/Mask wheel which has showed erratic behaviour recently. The intervention was sucesful, the S/M wheel improived but it is sitll not completely reliable and it requires slit/grism re-alignment after every spectroscopic acquisition. The S/M wheel init switch was also restored. Bottom line: SofI is operational with 1-2 min of extra overheads in spectroscopic modes (Dec 13, 2005).
  • SofI is operational with somewhat larger overheads for slit and grism alignment (Nov 7, 2005).
  • Due to a new failure of the grism wheel mechanism, spectroscopy and polarimetry are NOT available, until further notice. We appreciate your understanding. Affected visitor and service mode programs will be informed (Apr 5, 2005).
  • An intervention to fix the problems with the grism wheel and the objectiveis was carried out from Jan 27 to Feb 11, 2005. Unfortunately, the grism wheel problem reappeared a week later. We are currenly having a new intervention (20-31.03). The instrument is warm and it won't be available for observations until the beginning of Apr. (Mar 20, 2005).
  • All spectroscopic and polarimetric modes are temporarily decomissioned because of frequent grism wheel slippage. The modes will be restored during the Jan-Feb 2005 intervention. Contact the instrument scientist for details (Dec 2005).
  • The High Resolution Imaging Modes (SF, LF+FE and HR) are decommissioned as their image quality was unacceptable. The lenses are being fixed. These instrument capabilities will be restored during an intervention planned for Jan-Feb 2005 (Dec 2004).
  • Due to a new failure of the grism wheel mechanism, spectroscopy and polarimetry are NOT available, until further notice. We appreciate your understanding. Affected visitor and service mode programs will be informed (Apr 5, 2005).
  • An intervention to fix the problems with the grism wheel and the objectiveis was carried out from Jan 27 to Feb 11, 2005. Unfortunately, the grism wheel problem reappeared a week later. We are currenly having a new intervention (20-31.03). The instrument is warm and it won't be available for observations until the beginning of Apr. (Mar 20, 2005).
  • All spectroscopic and polarimetric modes are temporarily decomissioned because of frequent grism wheel slippage. The modes will be restored during the Jan-Feb 2005 intervention. Contact the instrument scientist for details (Dec 2005).
  • The High Resolution Imaging Modes (SF, LF+FE and HR) are decommissioned as their image quality was unacceptable. The lenses are being fixed. These instrument capabilities will be restored during an intervention planned for Jan-Feb 2005 (Dec 2004).

  • Send comments to: ls-sofi <ls-sofi[@]eso.org>
    This page created: June 13, 2005, by Valentin D. Ivanov. Last Update: Dec 13, 2006, by Valentin D. Ivanov
    [ESO][ESO][ESO][Search][Help][News]